OpenStudy (anonymous):

What is the equation of the line tangent to y=(x+3)^1/3 at x=-3? How do I find this? I am struggling. Thank you!

OpenStudy (anonymous):

dy/dx = ((x+3)^(-2/3))/3 therefore at x = -3 -3+3 = 0 = 0^n = 0 hence 1/0 = ND

OpenStudy (anonymous):

This is odd. However it should make sense considering y = 0 when x = -3...

OpenStudy (anonymous):

Ok. I find the critical points (0,-3) earlier but is there an equation? I've been doing these all morning and I'm thoroughly confused haha.

hartnn (hartnn):

so you got dy/dx at x=-3 as 0, right ?

hartnn (hartnn):

that means your slope of the line = m = 0

hartnn (hartnn):

you just need y- intercept when x= -3 what is the value of y ?

OpenStudy (anonymous):

@hartnn 0?...

hartnn (hartnn):

thats correct! so, in y = mx+c m=0, c=0 so, whats the equation ?

OpenStudy (anonymous):

@hartnn Do I plug those in to y=mx+c? y=0(x)+0

hartnn (hartnn):

yes

hartnn (hartnn):

y=0 x axis thats it! thats your equation of line :)

OpenStudy (anonymous):

Okay, thank you both SO much